Journalist Report – April 20th

Crew #: 335

Position: Crew Journalist

Report prepared by: Matteo Calore

Date: 21-04-2026

Sol: 1

First Day on Mars: Entering New Worlds

Today was our first day on Mars. We woke up around 6:30, and Charlotte made pancakes, which we happily devoured. Then we prepared for the EVA. Kat wasn’t feeling very well, so Elisa took her place for the first EVA to Marble Ritual between 10 and 11. Afterwards, Dave and Charlotte carried out the second EVA to Marble Ritual.

After a quick lunch of mashed potatoes and mushrooms prepared by Matteo, Dave, Elisa, and Matteo went out to film and work on photographic projects, while Kat and Charlotte stayed at the Hab to work and write. This evening Charlotte prepared a lentil and potato stew for dinner, which was very good, and everyone is now busy writing reports and planning the next days.

Arriving on Mars is an interesting experience. Despite months of online calls, we are only now beginning to really get to know one another, and to learn how to communicate our different needs. While some have known each other for a long time, others have only just met and come from very different worlds. We are slowly learning how to enter one another’s worlds.

Everyone is a bit tired and some feel sick. We are still adjusting from our different journeys, still partly carrying the baggage we brought from Earth, and trying to understand how this world works. It is a completely new reality. The desert around us is astonishing, although it also feels as if we are never really going outside, since we bring the Hab with us even when we go out, in our suits.

GreenHab Report – April 20th

Report title: GreenHab Report
Crew #: 335
Position: GreenHab officer
Report prepared by: Charlotte Brugge
Date: 20-04-2026
Sol: day 1

Environmental control (fan & heater): Automated
Average temperatures (last 24h): 86.9
Maximum temperature (last 24h): 105.9
Minimum temperature (last 24h): 67.9
Hours of supplemental light: Automated
Daily water usage for crops: 4.5 Liters
Daily water usage for research and/or other purposes: 0
Water in the Tank (160 gal useful capacity): 108.32
Time(s) of watering for crops: 13.30
Changes to crops: the small tomato seems more red, maybe tomorrow there will be a very small harvest
Narrative: watered and checked the plants.
Harvest: (include which crop and mass in grams) 0
Support/supplies needed:0

EVA Report – April 20th

Report title: EVA Report

Crew #: 335

Position: greenhab officer

Report prepared by: Charlotte Brugge

Date: (20-04-2026)

Sol: Day 1

EVA #: 2

Purpose of EVA: Marble ritual

Start time: 10.30

End time: 11.50

Narrative: we did decompression at 10.30, drove out to Marble ritual to walk and climb up a bit to look at the station. Dave took some pictures

Destination per MDRS Map:

Destination Coordinates (use UTM WGS 84): 4250800N, 518750W

EVA Participants: David Laude + Lot Brugge

Road(s) and routes per MDRS Map: The EVA will begin on Moonwalk Road, then proceed along Cow Dung Road 0110 to Marble Ritual.

Mode of travel: Rover + walk

Operations Report – April 20th

Report title: Operations Report

Crew #: 335

Position: Engineer

Report prepared by: Katherine Berry

Date: 20-04-2026

Sol: 1

Non-nominal systems: Radio 2 & 4. Headset 6.

Notes on non-nominal systems: Radio 2 exhibits intermittent connectivity at the female-end connector, with a stable connection only achieved when pressure is applied. Cross tested with other radio to confirm the fault is isolated to Radio 2, indicating a likely connector or internal contact issue.

Headset 6 has a nonfunctional push to talk switch, preventing transmission.

Radio 4 also demonstrates intermittent connectivity at the connector interface, requiring applied pressure/ manipulation for operation. Additionally, the unit has a partial communication fault: inbound audio is received, but outbound transmission is not detected by the Radio 4 operator.

ROVERS

Spirit rover used:

Hours: 306

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Opportunity rover used EVA 1:

Hours: 267.9

Beginning charge: 100%

Ending charge: 83%

Currently charging: Yes

Opportunity rover used EVA 2:

Hours: 268.0

Beginning charge: 100%

Ending charge: 47%

Currently charging: Yes

Curiosity rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used:

Hours: 405

Beginning charge: 100%

Ending charge: 43%

Currently charging: Yes

General notes on rovers: None

Summary of Hab operations:

Water Use (see notes): 31.01 gal/d

Main tank level (remaining gallons, see notes): 320

Main water tank pipe heater (ON or OFF): OFF

Main water tank heater (ON or OFF) OFF (was removed)

Toilet tank emptied (NO or YES): NO

Summary of internet: Connection was used for reading, research projects, accessing homework, sending reports, email and listening to music.

Summary of suits and radios:

David Laude, Commander: Suit 9, radio 3, headset 6

Elisa Strinna, Artist: Suit 8, radio 3, headset 1

Matteo Calore, HSO: Suit 9, Suit 7, radio 2, headset 2

Lot Brugge, Suit 10 , radio 4, headset 2

Summary of GreenHab operations: Watering

Water use: 1.2 gal

Heater (ON or OFF): OFF

Supplemental light (hours of operation): Auto

Harvest (name, weight in grams): 0g, N/A

Summary of ScienceDome operations: EVA note taking

Dual split (Heat or AC, On or Off or Automatic): Auto

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: Two crewmembers experiencing low grade fever, fatigue, headache and sore throat.

Questions, concerns and requests to Mission Support: We are unable to fix the electric can opener, would like to request a replacement if possible.

EVA Report – April 20th

Crew #: 335

Position: Crew commander

Report prepared by: Dave Laude

Date: 20-04-2026

Sol: 1

EVA #: 3

Narrative:

The purpose of this EVA was to survey much ground along travel route for film makers to consider for filming during the EVA or a later return. We will stop along the route as needed for filming. Also, for commander’s panaramic photo survey.

The EVA began at 1:30 PM heading north on Co Dung Rd. heading to final destination of Green Mars overlook, but Perseverance, with two passengers and heavy professional video camera and super duty tripod, charge dropped to 50% at Brahe Highway intersection. We decided too far to continue so turned around. During the trip there were four excellent filming stops thus making good use of the 3 hours. Returned at 4:30 PM. Dave’s panoramic photo project accomplished half of intended goals due to turning back. It will be reviewed to see if successful or if changes needed.
Destination : Green Mars View

Coordinates (use UTM WSG 84) :

Green Mars View : 4257000N, 520500W

EVA Participants:

Dave Laude (Commander)

Matteo Calore (HSO/Journalist)

Elisa Strinna (Artist)

Road(s) and routes per MDRS Map: The crew departed the MDRS via Moonwalk Road, then follow Cow Dung Road to Brahe Highway intersection

Mode of travel: driving

EVA Report – April 20th

Report title: EVA Report
Crew #: 335
Position: Crew Artist
Report prepared by: Elisa Strinna
Date: 19-04-2026
Sol: 1
EVA #: 1
Purpose of EVA:Practice EVA and look around for desiccated soil samples and for film making
Start time: 10:00
End time: 10:52
Narrative: We started preparation at 8:39 AM, we started the decompression at 8:55 AM, got in the rovers (9:00 AM ) and started our way to the destination. We arrived at 9:10 AM. We explored the area on the right side of the road for a little bit, took some pictures and short videos, and headed back at 10.38 AM. We arrived to the station at 11.46 AM. We started decompression and entered the hub at 11.52
Destination per MDRS Map: Marble Ritual
Destination Coordinates (use UTM WGS 84): 4250800N, 518750W
EVA Participants: Matteo Calore and Elisa Strinna
Road(s) and routes per MDRS Map: The EVA will begin on Moonwalk Road, then proceed along Cow Dung Road 0110 to Marble Ritual.
Mode of travel: Spirt

Sol Summary – April 20th

Report title: Sol Summary Report

Crew #: 335

Position: Crew Commander

Report prepared by: Dave Laude

Date: 20-04–2026

Sol: 1

Summary Title: Three EVAs on day one.

Mission Status: Going fair

Sol Activity Summary:

We arrived on Mars midnight and awoke to views of Mars out the portals. Fantastic views! Charlotte made pancakes for breakfast. Who doesn’t like them? EVAs 1 and 2 took place in the morning to Pebble Ritual as in normal practice. One of us was not feeling up to it and didn’t EVA. Afternoon the film makers and commander headed to Green Mars Overlook, however once reaching the Brahe Highway intersection rover Perseverance battery was at 50% with significant distance remaining ahead and so we turned back. However, we made full use of 3 hours with much filming along the way. Another crew member is now also not feeling well now with sore throat. Our HSO monitors both.

Look Ahead

Tomorrow two EVAs, this time focusing on Charlotte’s clay project while film makers film her.

Anomalies in work:

Once reaching the Brahe Highway intersection, rover Perseverance battery was at 50% with significant distance remaining ahead and so we turned back.

Weather:

It was sunny today, up to 80F and very mild winds.

Crew Physical Status:

Two show mild signs of virus.

EVA:

We ran a little low on battery, but otherwise very good.

Reports to be filed:

Sol Summary
Operations Report
Greenhab Report
Journalist Report
EVA Report
EVA Requests
Daily photos
Support Requested:

Supplemental Operations Report – April 17th

Date: 4/17/2026
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Hab outer shell. GreenHab hydrant. Director’s trailer roof.

Power system: "1) Power system: Nominal.
2) Current generator hours: 1448.0"

Propane Readings:
Station Tank: 49%
Director Tank: 60%
Intern Tank: 70%
Generator Tank: 80%

Water:
Hab Static Tank – 416 gallons above safety margin
GreenHab – 111 gallons above safety margin
Outpost tank – 400 gallons
Septic Tanks emptied: No

Rovers:
Sojourner rover used: Yes, around campus
Hours: 226.9
Beginning Charge: 100%
Ending Charge: 100%
Currently Charging: Yes
Opportunity and Perseverance charged fuse box replaced due to their malfunction. A new SOP is created for this type of repair. Curiosity driver seat replaced.

Cars:
Hab Car used and why, where: Multiple trips to Hanksville for supplies
Crew Car used and why, where: Replacement vehicle is still being sourced
General notes and comments: The station will be operating with the HAB car only until crew car function is replaced

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal

Campus wide inspection, if action taken, what and why: HAB view camera installed on top of RAM. SciDome driveway and RAM view cameras replaced. SciDome junction tarp replaced and reinforced with ropes. GreenHab junction tarp reinforced with ropes. Hab outer shell shows ongoing wear.
Summary of Hab Operations: Nominal. L-Deck main and porch lights remote control relay replaced. L-Deck EVA room, main airlock and restroom lights remote control relay installed. SIMOC sensors reprogrammed to be connected to our Dashboard. Emergency signal lights installed for testing. Toilet seal replaced. HVAC air filter replaced.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal. Main power meter got off-line, it was rebooted by cycling breaker 3 and 5 simultaneously.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Nominal.
Summary of Outpost Operations: Intern trailer: thermostat replaced to enable A/C function; Zigbee-wifi gateway installed to enable zigbee IoT at the Outpost; Atm sensor installed; main heater DC motor replaced. Director trailer: Atm sensor installed; roof is leaking around A/C and along the eastern side of the trailer. Shed lights remote control relay installed.
Summary of Health and Safety Issues: Nominal.

Mission Summary – April 10th

Report title: Summary mission Report
Crew #: 333
Position: Crew Commander
Report prepared by: Marie Jansen
Date: 10-04-2026
Sol: SOL 12

Summary Title: Wrapping things up
Over the course of the twelve Sols, Crew Arsia established a rhythm that became the heartbeat of the mission. From the very first days, the team settled into a routine built around shared meals, scientific work, EVA operations, and the quiet moments in between that slowly turned six individuals into a cohesive crew.
Mornings typically began at 7:00 or 8:00 AM, almost always anchored by a communal breakfast. Pancakes became something of a crew tradition, appearing on the menu more often than any other dish, though porridge, granola with smoothies, and cereal also made regular appearances. These morning meals were never just about food, they were the moment the crew came together before splitting into EVA teams and hab teams for the day ahead.
EVAs were the backbone of the mission. The crew ventured out regularly to carry out their scientific experiments across a variety of sites around the MDRS, including Cowboy Corner, Green Mars View, Kissing Camel Ridge, Candor Chasma, Somerville Overview, North Ridge, and the Special Region. Each location brought its own character, some humbling in their vastness, others charming in their unexpected beauty. Equipment challenges were a recurring theme: rovers that failed to recharge properly, a Tap Strap that ran out of battery at a critical moment, microphone issues, and communication difficulties with CapCom in the field. Each time, the crew adapted, reorganized, and pushed forward without losing focus on the mission objectives.
Back at the hab, those not heading out kept equally busy. The GreenHab required regular attention, and Matthias took on much of that responsibility throughout the mission: tending to the plants, carrying out his seed analysis, and keeping a close eye on the greenery that sustained both morale and science. A running mystery added some unexpected excitement when the crew discovered that an animal had been sneaking into the GreenHab and nibbling on their plants. After several hours of suspicion, the culprit was eventually caught, much to the satisfaction of the entire team.
Meals were a cornerstone of daily life. The crew made a point of putting together hearty, nourishing dishes to keep everyone fueled through long days of physical and mental effort. Highlights included Zahraa’s couscous, Matias’s omelettes and tuna burgers, a memorable burger night, chicken curry, mac and cheese, cheddar nachos, and special Easter muffins prepared by Joanna and Marie. Cooking became an act of care, and the dinner table was always a place to reconnect at the end of the day.
The mission was also an emotional experience. A rest day was dedicated to celebrating Easter together, with a chocolate chip pancake breakfast, an egg hunt in the hab, and an afternoon photoshoot. There were card games, movie nights, mimes, and cycling sessions on the stationary bike. There were also harder moments, tensions within the group surfaced at times, leading to an open team meeting that helped bring things back into balance. It was a reminder that the human side of space exploration is just as demanding as the technical one.
As the Sols passed, the crew completed experiment after experiment, explored corner after corner of the surrounding terrain, and slowly worked their way toward the finish line. By the final Sol, with bags being packed and journals being filled, one thing was clear: Crew Arsia had not just conducted a simulation. They had lived something real.

Sol Summary – April 10th

Report title: Sol Summary Report
Crew #: 333
Position: Crew Commander
Report prepared by: Marie Jansen
Date: 10-04-2026
Sol: SOL 12

Summary Title: Wrapping things up
Mission Status: last day…
Sol Activity Summary:
And so we find ourselves on the final day of our simulation. We had planned to enjoy a slow, leisurely morning, but a drone had other ideas, buzzing over the base and waking the crew up, so the day ended up starting at 8:00 AM after all. We made the most of it with a batch of pancakes to kick things off on a good note.
The morning was then dedicated to wrapping up the last remaining experiments before the simulation came to a close. Zahraa, Antoine, and Joanna each completed their final runs, and Matthias is scheduled to carry out his seed analysis later this afternoon.
There is a quiet sadness hanging over the hab today as the reality of the ending sinks in. But above all, the crew is filled with an immense sense of pride and gratitude for everything we have experienced and built together over these past days
Look Ahead Plan:
Nothing is planned tomorrow on the mission front, we will carry out a final walkthrough of the station, hand things back as we found them, and say goodbye to the MDRS. Thank you for everything.
Anomalies in work: none
Weather:
A perfect day to bring the simulation to a close.
Crew Physical Status:
Everyone is doing well! We all made it out alive!
EVA: None
Reports to be filed:
Sol Summary
Operations Report
Greenhab Report
Journalist Report
Astronomy Report
EVA Report
Daily photos
Mission summary
End-mission research report
Food inventory
Support Requested: none